OpenStudy (anonymous):

Solve the system of equations below by graphing both equations with a pencil and paper. What is the solution? y = -x - 1 y = x + 3 A. (0, 3) B. (-1, 2) C. (-2, 1) D. (0, -1)

OpenStudy (anonymous):

solve \[x+3=-x-1\] in three steps a) add x to both sides to get \(2x+3=-1\) b) subtract 3 from both isdes to get \(2x=-4\) c) divide by 2 to get \(x=-2\)

OpenStudy (anonymous):

so as usual the answer is C it is always C